Demonstrate that the hash function
H(x) = (7x + 13) mod 31
is not weakly collision resistant, for H(5), by showing how easy it is to find such a collision.
please show your work!!
Answer #1

Answer 8- consider the hash function HC = ( 7a+ 13) mod 31 H15) =(745 +13) mod 31 = 48 mod 31 = 17. - Thus, the value of H (5